### v1

# In-context Learning for Automated Driving Scenarios

Abstract
One of the key challenges in current Reinforcement Learning (RL)-based Automated Driving (AD) agents is achieving flexible, precise, and human-like behavior cost-effectively. This paper introduces an innovative approach utilizing Large Language Models (LLMs) to intuitively and effectively optimize RL reward functions in a human-centric way. We developed a framework where instructions and dynamic environment descriptions are input into the LLM. The LLM then utilizes this information to assist in generating rewards, thereby steering the behavior of RL agents towards patterns that more closely resemble human driving. The experimental results demonstrate that this approach not only makes RL agents more anthropomorphic but also reaches better performance. Additionally, various strategies for reward-proxy and reward-shaping are investigated, revealing the significant impact of prompt design on shaping an AD vehicle’s behavior. These findings offer a promising direction for the development of more advanced and human-like automated driving systems. Our experimental data and source code can be found here [^1] GitHub: https://github.com/JingYue2000/In-context_Learning_for_Automated_Driving..
Index Terms:
Large language models, Automated driving, Reinforcement learning, DQN, In-context learning
I Introduction
With the advancement of artificial intelligence, an increasing number of machine learning techniques are being applied in the field of intelligent transportation [1][2]. Among them, Reinforcement Learning (RL) is a subfield of machine learning that deals with how an agent can learn to make decisions and take actions in an environment to achieve specific goals following a trial-and-error learning approach [3]. Reinforcement Learning showcases robust learning and adaptability, particularly in real-time decision-making scenarios prevalent in automated driving [4, 5, 6].
Despite its potential benefits in enhancing safety [7], efficient navigation, and reducing traffic congestion in real-world driving [8], there are challenges hindering its widespread adoption [9]. For example, one notable challenge is designing accurate reward functions that encourage safe, efficient, and human-like driving behavior; yet, there are other difficult aspects to consider as well:
  • Perfecting reinforcement learning agents to mirror human driving while ensuring safety and efficiency remains a complex task.
  • Creating fair and easy-to-understand reward functions remains challenging due to driving task and environment complexity.
  • Traditional methods, heavily reliant on extensive labeled data for designing reward functions, often fail to adapt to variable user needs or evolving objectives in new contexts.
To solve these challenges, a shift in approach is emerging with In-context learning (ICL),  [10], where LLMs make predictions only based on contexts augmented with a few examples [11]. By specifying an objective with a natural language prompt, LLM can convert user goals into reward signals, and guide RL agents intuitively toward desired outcomes [12]. This innovative interaction between LLMs and RL agents not only advances opportunities in fields like automated driving but also highlights the LLMs’ versatility in complex decision-making.
Building on this innovation, we introduce our approach for leveraging the synergy between LLMs and RL agents in the context of automated driving scenarios. We explore the mechanisms by which the reward guidance offered by LLMs influences the training and performance of RL agents. Further, we explore methodologies to achieve diverse driving behaviors through the integrated LLM-RL framework, demonstrating the potential of this collaboration in enhancing autonomated driving technologies.
The contributions of this paper are summarized as follows:
  • We proposed an RL training framework incorporating LLMs to simplify RL-AD agents’ reward design in AD scenarios, quoted as LLM-RL agent in this paper.
  • We explore the integration of LLMs with RL agents to enable different human-like driving styles, allowing for dynamic adaptation.
  • We conducted a comparative analysis of previous approaches and showed that the integration of LLMs into the RL training process leads to more objective-aligned learning outcomes.
The remainder of this paper is organized as follows. Section II provides an overview of the related work followed by presenting the problem formulation and approach in III. The proposed approach is further described in Section IV. The experiments and results are discussed in Section V. Finally, the conclusion and future work are summarized in Section VI.
II Related Work
II-A Large Language Models in Reinforcement Learning
Large Language Models (LLMs), known for their capacity to learn from minimal context and deliver human-like responses following training with human data [13], are increasingly considered as a potential solution when combined with Reinforcement Learning (RL) agents to perform higher-level tasks. In certain studies, RL is employed to fine-tune existing LLMs[14, 15, 16]. Conversely, other research integrates RL with LLMs to execute specific tasks or develop particular skills[17, 18, 19]. However, our research diverges from these methods by using LLMs as proxy rewards, enhancing RL model training as described in[20].
The research landscape in using large language models (LLMs) as proxies for reinforcement learning (RL) rewards has varied approaches. Kwon et al. pioneered the use of LLMs for learning in contexts such as the Ultimatum Game and negotiation tasks, focusing solely on LLM-generated rewards. [12]. In contrast, our study utilizes a diverse set of rewards in dynamic AD system scenarios, broadening the scope of possibilities. Other researchers, like Xie et al. and Ma et al. [21][22], have explored reward design through direct Python code generation. Meanwhile, Song et al. in [23] required human inputs to iteratively fine-tune rewards, whereas another study [24] used LLM outputs as auxiliary rewards. Unlike these approaches, our method directly employs LLM responses to generate reward signals, simplifying the process and eliminating the need for human intervention.
II-B Large Language Models in Automated Driving
Utilizing Large Language Models (LLMs) in automated driving enhances decision-making by tapping into LLMs’ vast knowledge and reasoning capabilities. Research in vehical planning and control mainly falls in performing Prompt Engineering  [25, 26]or Fine-tuning Pre-trained Model as described in  [27, 28].
Targeting at solving closed-loop driving tasks, Fu et al.  [26] utilize LLMs to comprehend and interact with environments. Wen et al. improved this framework and help LLM make decision with different reflection modules. [25]. Our approach simplifies decision-making by considering only the ego car’s most recent action. Similarly, Cui et al.  [29] proposed a method that can provide the LLM with comprehensive environment information from data collected from different modules of the vehicles, ensuring safer results from LLM decisions. Despite these advancements, these approaches suffer from real-time latency in decision-making.
Other researchers focused on fine-tuning the pre-trained model, like Chen et al.  [28] performed fine-tuning with 10,000 generated driving scenarios. DriveMLM [30] merged linguistic processing with vehicle control for a more advanced system. Our work leverages LLM to enhance RL performance, satisfying real-time requirements by dynamically querying it based on scenarios.
III Problem Formulation and Modeling
III-A Markov Decision Process Formalization
An automated driving scenario can be modeled as a Markov Decision Process: $M=\langle S,A,\rho,R,\gamma\rangle$, where $S$ is the state space representing various driving scenarios (e.g., traffic density, relative speeds of vehicles, lane positions, etc.). $A$ is the action space representing the set of all possible driving actions (e.g., accelerate, decelerate, change lanes, etc.). $\rho:S\times A\times S\rightarrow[0,1]$ is the transition probability function representing the likelihood of transitioning from one state to another given an action. $R:S\times A\rightarrow\mathbb{R}$ is the traditional reward function mapping states and actions to real-valued rewards. $\gamma$ is the discount factor, indicating the agent’s consideration for future rewards.
Inspired by [12] we employed a LLM as the proxy for reward calculations in our framework to achieve a more interpretable and adaptable reward mechanism.
Figure 1: LLM acts as a reward proxy.
As illustrated in Fig. 1 our integrated framework uses the LLM to interpret textual prompts $\varrho$ as input and outputs a reward signal. The text prompt $\varrho$ consists of three major components: Task Description $\rho 1$, User Objective $\rho 2$ and Last Outcome $\rho 3$. The LLM is modeled as a function LLM: $A*\rightarrow A*$, taking the concatenated prompt $\varrho$ as input and outputting a string. A parser $A*\rightarrow{0,1}$ is then used to map the textual output of the LLM to a binary reward signal, which refers to the (2) of Fig. 1. Then, the binary reward signal will be sent to the RL agent for further training, which refers to the (3) of Fig. 1. This framework replaces the traditional reward function $R$ with a proxy reward function LLM and can be used with any RL training algorithm.
To explore whether LLM-assisted agents could achieve a more balanced and safe driving behavior, emulating human-like decision-making in these scenarios, we implemented our framework in highway scenario as explained in Section IV.
III-B Reward Function Shaping Strategies
LLM’s response subsequently translated into a quantifiable reward signal. This reward, termed "LLM Reward" in this paper, is an integer, which is 0 or 1, reflecting the appropriateness of the agent’s actions as evaluated by the LLM. This innovative approach aims to leverage the nuanced understanding capabilities of LLMs to refine and guide the RL agent’s decision-making process.
We used combined reward to take advantage of both Deep Q Learning (DQN) and LLM, referring to the method mentioned in [31]. Given the original reward function in a highway driving environment, we propose an adjusted reward function that integrates a reward signal from LLM and scales the combined reward to the [0,1] range using a tangent function:
$R_{\text{ad}}(s,a)=\sigma\left(\text{Env\_reward}+\cdot\text{LLM\_reward}(s,a)\right)$ (1)
[…]
IV Proposed Approach
A simulation environment customized by HighwayEnv [32] is used. HighwayEnv is capable of mimicking real-world driving scenarios to generate scene descriptions. We selected the GPT-3.5-turbo-1106 model from OpenAI for its robust performance, fast response times, and suitability for handling the extensive data involved in our research.
A standard DQN model is set up as a baseline agent, trained with reward function 2 that are calculated based on the current driving scenario, which serves as a foundation of our proposed approach.
In the proposed approach, a prompt for one interaction consists of the last action, the predefined prompt, and the scene description. This description includes available actions, a general overview, and a safety evaluation. As shown in Fig. 2, each arrow directed towards the agent symbolizes a segment of the prompt, sourced from the categories previously mentioned, together forming a comprehensive and effective prompt.
Figure 2: Prompts structure
These prompts distinctly outline the scenario as well as instruction for LLM to make decision for further steps.
Scenario informations included are: speed of the vehicle in control(the LLM-RL agent), speed and positions of nearby vehicles, LLM-RL agent’s last action and possible actions to choose from.
Further instructions including driving rules, caution points and specified format needed for generating and presenting responses.
IV-A Basic Driving Rules and Task Description
To illustrate the context for the ChatGPT agent, we provide a series of predefined prompts:
  • SYSTEM_MESSAGE: Specifies LLM’s role and setting the scenario’s context.
  • Driving_RULES: Provides the general safe-driving guidelines which shapes the behaviour pattern for the driving styles of the agents.
  • DECISION_CAUTIONS: Explains how to make well-informed decisions.
These predefined messages and rules can serve as a foundation for the behaviour of the driving assistant. The construction and phrasing of previous descriptions are based on the work of drive-like-human [33].
IV-B Context Description
To illustrate the context of the current road situation, the following basic elements are defined as Lane and Vehicle. The Vehicle class encapsulates properties including speed in both directions and positional data, while the Lane class incorporates attributes including position, unique identification, and details regarding adjacent lanes.
IV-C Observation Converting and Integrating Language Model
To integrate Language Model-driven insights into the Reinforcement Learning framework, we converted the RL agent’s current observation and actions into a natural language context through several analysis functions, which could be processed by LLM. This approach ensures that each response from the LLM is based on both current input and past action.
IV-D Formatting Output
This part of the prompt aims at formatting an output string that is interpretable to our program during runtime. We specify the outputs of the agent to generate only the following five actions of the ego vehicle: LANE_LEFT, IDLE, LANE_RIGHT, FASTER, and SLOWER. The LLM interprets all information above context and generates feedback. A procedure to generate a final reward is discussed as follows:
We initialize Highway_Env and input a concatenated prompt, which includes the last action, the pre-define prompt, and the scene description. The last action is derived from the previous decision made by the agent. This prompt enables the agent to understand the scenario it faces and ensures that the responses generated by the agent adhere to user requirements.
By comparing the answer provided by the LLM with the one provided by the RL algorithm, a binary LLM reward is generated. If the decisions match, the reward is 1; otherwise is 0, refereed to LLM reward in this paper. Simultaneously, the RL agent obtains its reward in the given environment. These rewards are then weighted and summed, using the tanh function for normalization to ensure the final reward falls within the range of 0 to 1.
V Experiments and Analysis
V-A Highway AD as a case study
The practical capabilities of leveraging the in-context learning ability of LLM in AD are illustrated through highway driving scenarios. The potential for AD agent personalization in highway scenarios by establishing two groups with different meta-prompts is also explored in this case study. In the prompt case shown in Fig. 3, the scenario can be described as the ego car driving in lane 1 on a highway. There is one vehicle in its current lane. The right lane change is deemed safe by the safety assessment. The previous decision the ego car made, was to stay idle, which means the ego car is in the process of following the front car. This scenario featured two distinct driving agents: the "aggressive" and "conservative" agents. A comprehensive analysis was conducted under controlled conditions to compare their performance.
Figure 3: Meta prompt case study
Each RL agent’s driving behaviour was shaped through a series of user instructions, establishing their respective driving style within the scenario as shown in Fig. 4. Agent’s decisions show strong consistency in humanized behavior reflecting the behavior pattern meta-prompt.
The agent consider several factors: Safety, Efficiency and Driving Rules. For an aggressive agent, the user has stated that efficiency is of higher priority, hence the agent is considering lane right for higher efficiency. For a conservative agent, the user has stated that safety is of higher priority, hence the agent is considering slowing down to keep distance from the leading car.
Figure 4: RL agent’s driving behaviour
Compared with manually designing a reward function, this manner is simpler and has demonstrated its effectiveness in guiding the behaviour of an agent. These observations indicate that Large language models (LLMs), such as GPT in our case, offer a potential cost-efficiency alternative to performing similar roles with human feedback in certain contexts. LLMs can leverage vast amounts of text data to learn implicitly from human-generated content, interactions, and feedback, significantly reducing the human workload in reward design while achieving similar effects.
V-B RL Agent Selection and Implementation
DQN was chosen for our experimentation as it’s rather efficient and satisfied our research requirement. It shows suitability in handling discrete action spaces like AD problems and thus could be utilized to assess whether LLM could effectively guide RL in our scenario.
V-C Experiment Design and Performance Evaluation
Mean reward, lane change percentage and speed up percentage are used as metrics to evaluate the performance of our agent. Mean reward represents the average total reward for actions taken in the testing scenario. Lane change percentage indicates the proportion of lane-changing actions out of the total actions, while speed-up percentage represents the proportion of acceleration actions out of the total actions. The agents are first trained in a customized environment. After training various agents, we conducted multiple experiments and analyses to assess their performance within a consistent and standardized highway testing environment.
Does LLM reward guide the training process of RL agents?
Using the sum of LLM reward and RL reward as a combined reward function for reinforcement learning and comparing it with the DQN baseline. We compare the performance by testing the average reward function obtained by running the trained agent in the same environment for 5 rounds. The results are as follows: The match rate increases through training aligning with the guidance of LLM. The match rate remains steady at around 50 per cent as shown in Fig. 5.
Figure 5: Match rate through training
Does prompt affect the performance or behaviour pattern of the LLM-RL agent?
Three agents were designed and implemented with different driving styles. The base driving agent, which utilizes the LLM’s general knowledge of driving without customized prompts, serves as the foundation for specialized adaptations. Specifically, the aggressive driving agent prioritizes efficiency, whereas the conservative driving agent emphasizes safety. The experiments were conducted in the same highway environment. Compared to the DQN baseline, both the conservative and base agents demonstrate relatively low lane-change and speed-up rates as shown in Fig. 6. This confirms that the behavioral patterns of these agents diverge from those of DQN agents, which frequently rush and make frequent lane changes to access higher rewards. Specifically, the aggressive agent displays a notably high speed-up rate compared to all agents, showcasing the influential effect of prompts.
(a) Collision Probabilities for different driving style agents
(b) Average speed for different driving style agents
Figure 6: Training Results
Behaviour analysis in terms of mean reward, lane change and speed up are summarized in table I.
TABLE I: Experiment result: behaviour analysis
Index Mean Reward Lane Change Speed Up
DQN baseline 0.82824 0.30681 0.42045
Aggressive 0.83888 0.02326 0.83721
Conservative 0.71391 0.01333 0.00666
Base 0.80140 0.10345 0.10345
Detailed rewards are shown in Table IIbelow, confirming the previous conclusion. An aggressive agent denotes a rather high-speed reward compared to all other agents. Specifically, the highway-env reward can be broken down into four rewards in different aspects: collision reward, right lane reward, high-speed reward, and on road reward.
TABLE II: Experiment result: reward break down
Method Collision Lane Change High Speed On Road
Reward Reward Reward Reward
DQN baseline 0.02632 0.51316 0.29933 1.0
Aggressive 0.14706 0.41177 0.69912 1.0
Conservative 0.0 0.5 0.01435 1.0
Base 0.02255 0.58271 0.17517 1.0
The behaviour patterns of different agents are shown in the videos available at[34]. A conservative agent denotes a more stable behaviour and prioritises safety, which can be investigated by the 0 in collision reward. However, it also causes a lower speed referred to lower high speed reward. An aggressive agent shows a behavior pattern that prioritises efficiency. In this test case, the ego car takes frequent lane changes to maximize efficiency, to a certain extent, disregarding safety.
VI Conclusion
The conducted experiments with LLM-RL integration showcased promising directions for incorporating language models into reinforcement learning. We proposed the LLM-RL AD framework, which includes an in-time query for LLM reward and utilizes a combined reward function. Extensive experimental results showcase LLM-RL’s capability to perform customizable behavior and demonstrate LLM can guide RL agents to achieve superior performance in complex environments like HighwayEnv. Moreover, our exploration of prompt design also shows that precise LLM responses are better than multiple vague options. Though the LLM’s reward can guide the training process of RL agents and prompts can affect their behavior patterns, the framework might not fully address mistakes made by the LLM, and there remains a need for the LLM to develop a deeper understanding of its surroundings to ensure accurate decision-making.
Future improvements could capitalize on finding more proper prompts and effective driving rules, aiming to help the LLM comprehend the scenarios better, which leads to more efficiency. For example, defensive driving scenarios can be considered based on Responsibility-Sensitive Safety (RSS) approach [35]. RSS highlights five safety rules an automated driving vehicle should be able to follow. These five rules include safe distance (i.e., the self-driving vehicle should not hit the vehicle in front), cutting in (the self-driving vehicle should be able to identify when lateral safety may be compromised by a driver unsafely crossing into its lane), right of way (the self-driving vehicle should be able to protect itself against human drivers who do not properly adhere to the right of way rules), limited visibility (be cautious in areas with limited visibility), and avoiding crashes (the self-driving vehicle should avoid a crash without causing another one).